The area A of a circle is increasing at a constant rate of 2cm2s−1. If the area of the circle is given by A = πr2, what is the rate of change of the radius when the radius is 4cm?


1
Expert's answer
2021-12-08T14:03:00-0500
"A=\pi r^2"

Differentiate both sides with respect to "t"

"\dfrac{dA}{dt}=\dfrac{d}{dt}(\pi r^2)"

"\dfrac{dA}{dt}=2\pi r\dfrac{dr}{dt}"

Solve for "\dfrac{dr}{dt}"

"\dfrac{dr}{dt}=\dfrac{\dfrac{dA}{dt}}{2\pi r}"

Given "\dfrac{dA}{dt}=2 {cm}^2/s, r=4cm"



"\dfrac{dr}{dt}|_{r=4}=\dfrac{2{cm}^2/s}{2\pi (4cm)}=\dfrac{1}{8\pi}cm/s\approx0.040cm/s"

The rate of change of the radius is "\dfrac{1}{8\pi}cm/s\approx0.040cm/s" when the radius is 4cm.